# On Geometric Evolution and Microlocal Regularity of the Navier–Stokes Equations

### v2

@@ V.1 Energy Functionals
Metric compatibility and control of geometric defects
A subtle but important feature of the present development is that the effective connection $\nabla^{u}$ introduced in Section III is, in general, not compatible with the effective metric $g^{\prime}$. More precisely, the non-metricity tensor
$(\nabla^{u}g^{\prime})_{ijk}:=\partial_{i}g^{\prime}_{jk}-(\Gamma^{u})^{l}{}_{ij}g^{\prime}_{lk}-(\Gamma^{u})^{l}{}_{ik}g^{\prime}_{jl}$ (106)
does not vanish identically unless additional structural conditions on the velocity field are imposed.
Recalling that $(\Gamma^{u})^{k}{}_{ij}=\Gamma^{k}{}_{ij}+A^{k}{}_{ij}$, we compute
$(\nabla^{u}g^{\prime})_{ijk}=-\bigl(A^{l}{}_{ij}g^{\prime}_{lk}+A^{l}{}_{ik}g^{\prime}_{jl}\bigr)$ (107)
which shows that the failure of metric compatibility is entirely encoded in the deformation tensor $A(u)$ associated with the velocity field. In particular, $\nabla^{u}$ is metric-compatible with respect to $g^{\prime}$ if and only if
$A^{l}{}_{ij}g^{\prime}_{lk}+A^{l}{}_{ik}g^{\prime}_{jl}=0$ (108)
a condition that is satisfied, for instance, when $u$ is covariantly constant.
Although $\nabla^{u}$ is not metric-compatible in general, this defect does not obstruct the energy method developed in this section. Indeed, since $A(u)$ is linear in $\nabla u$, the non-metricity tensor satisfies the pointwise bound
$\|\nabla^{u}g^{\prime}\|_{L^{\infty}(S^{*}M)}\leq\sigma\,\|\nabla u\|_{L^{\infty}(M)}$ (109)
with $\sigma$ depending only on the background geometry. As a consequence, all error terms generated by the lack of metric compatibility enter the energy balance as lower-order contributions and are absorbed into the constants appearing in the Grönwall inequality. This shows that the geometric dissipation mechanism remains robust despite the presence of controlled non-metricity.
Reconstruction of the physical solution
The microlocal formulation developed above must ultimately be compatible with the reconstruction of physical solutions on the base manifold $M$. For scalar quantities, the reconstruction is performed fiberwise by
$\phi(x,t)=\int_{S_{x}^{*}M}\tilde{\phi}(x,t,\xi)\,d\xi$ (110)
where $d\xi$ denotes the canonical measure on the unit cotangent sphere at $x$.
Since $S_{x}^{*}M$ is compact and the microlocal lift $\tilde{\phi}$ is smooth in $(x,\xi)$ under the assumptions of Section V, the integral (110) is well-defined pointwise for all $(x,t)$. Moreover, by Hölder’s inequality and the definition of the microlocal functional $\mathcal{V}_{\phi}$, one has
$\|\phi(x,t)\|\leq\int_{S_{x}^{*}M}\|\tilde{\phi}(x,t,\xi)\|\,d\xi\leq\sigma\,\mathcal{V}_{\phi}(t)$ (111)
with $\sigma$ depending only on the dimension.
Differentiation under the integral sign is justified by the smooth dependence of $\tilde{\phi}$ on $x$, which yields
$\|\nabla^{k}\phi(\cdot,t)\|_{L^{\infty}(M)}\leq\sigma_{k}\sup_{(x,\xi)\in S^{*}M}\|\nabla_{x}^{k}\tilde{\phi}(x,t,\xi)\|$ (112)
for all integers $k\geq 0$. Consequently, uniform microlocal bounds on $S^{*}M$ propagate directly to uniform regularity bounds for the reconstructed physical solution. This shows that the microlocal framework is not merely auxiliary, but faithfully encodes the regularity properties of the original Navier–Stokes dynamics.
Geometric classification of admissible blow-up mechanisms and control of vortex stretching
@@
We now specialize to the vorticity $\tilde{\omega}$, whose geometric structure allows for finer control. Since $\tilde{\omega}$ is a closed microlocal $2$-form, we introduce the functional
$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\,\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}+\log\|\tilde{\omega}\|\right)\mu_{S^{*}M}$ (113)
@@
where $d^{\perp}$ denotes the vertical exterior derivative induced by the horizontal–vertical splitting of $TS^{*}M$, and $\tau>0$ is a dimensional parameter. This functional penalizes sharp angular gradients of vorticity and quantifies directional irregularity.
This functional $W[\tilde{\omega}$] admits a natural interpretation as a microlocal action measuring the angular regularity of vorticity on the compact phase space $S^{*}M$. The logarithmic term $\log\|\tilde{\omega}\|$ plays the role of an entropic density associated with the distribution of vorticity along cotangent directions, while the vertical gradient term $\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}$ penalizes sharp directional concentration on the fibers $S_{x}^{*}M$. From a dynamical viewpoint, the lifted flow generated by $\partial_{t}+\mathcal{L}_{X_{u}}$ defines a family of microlocal trajectories whose evolution is governed by a non-metric, torsion-bearing effective connection induced by the velocity field. The commutator structure developed throughout the paper shows that curvature, torsion, and non-metricity enter the dynamics as lower-order geometric forces. In this sense, $W$ can be viewed as an effective action functional on the space of microlocal paths, whose monotonicity expresses the suppression of highly curved or torsion-dominated angular trajectories by viscous dissipation. Persistent directional concentration would therefore require a breakdown of this geometric diffusion mechanism, linking potential singular behavior to a loss of microlocal dissipative stability rather than to amplitude growth alone.
A direct differentiation in time, combined with integration by parts along the fibers [30] and the compactness of $S^{*}M$, yields the dissipation inequality
@@
$\frac{d}{dt}W[\tilde{\omega}]\leq\int_{S^{*}M}\left(\tau\,\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}-\tau\,\|S\|\,\|\tilde{\omega}\|-\nu\right)\mu_{S^{*}M}$ (114)
@@
Since $S^{*}M$ is compact, the estimates derived above yield uniform control of the microlocal vorticity $\tilde{\omega}$ on $S^{*}M$. In particular, as long as the assumptions of Section V remain valid, no unbounded growth of $\tilde{\omega}$ can occur along the lifted dynamics. By the reconstruction formula, this microlocal control propagates to the physical vorticity field on $M$ [4]. Consequently, any loss of regularity of the Navier–Stokes solution must be accompanied by a breakdown of at least one of the hypotheses employed in the microlocal framework.
This observation highlights one of the central difficulties in the analysis of three-dimensional Navier–Stokes flows: the vortex stretching mechanism, which is responsible for potential amplification of vorticity. Recalling the classical decomposition of the velocity gradient, $\nabla u=S+\Omega$ into its symmetric and antisymmetric parts, only the symmetric tensor $S$ contributes to stretching effects.
Within the microlocal formulation, the evolution of $\tilde{\omega}$ along integral curves of the lifted transport field $X_{u}$ satisfies the pointwise differential inequality
$\frac{d}{dt}\|\tilde{\omega}\|\leq\|S(t)\|_{L^{\infty}}\,\|\tilde{\omega}\|-\nu\,\|\tilde{\omega}\|$ (115)
which makes explicit the competition between stretching and viscous dissipation. Integrating this inequality yields
$\|\tilde{\omega}(t)\|\leq\|\tilde{\omega}(t_{0})\|\exp\!\left[\int_{t_{0}}^{t}\bigl(\|S(s)\|_{L^{\infty}}-\nu\bigr)\,ds\right]$ (116)
This bound remains meaningful even in the vanishing viscosity limit $\nu\to 0$, in which case any amplification of vorticity necessarily requires uncontrolled growth of the time integral of $\|S\|_{L^{\infty}}$.
The vortex stretching mechanism is further constrained by the monotone microlocal volume functional $W[\tilde{\omega}$] introduced earlier. Since $S^{*}M$ is compact, the dissipation inequality satisfied by $W[\tilde{\omega}$] implies that sharp angular concentration of vorticity in phase space cannot persist unless accompanied by a loss of temporal integrability of the symmetric deformation, namely
$\int_{t_{0}}^{t}\|S(s)\|_{L^{\infty}}\,ds=\infty$ (117)
In particular, extreme directional focusing of vorticity in the cosphere bundle is geometrically incompatible with sustained dissipation. This shows that vortex stretching, while necessary for potential singular behavior, is subject to strong geometric and microlocal constraints imposed by viscosity, compactness, and the structure of the lifted dynamics. We therefore obtain a geometric exclusion principle: any blow-up scenario must simultaneously overcome viscous damping, geometric diffusion on the cosphere bundle, and the monotonicity of microlocal volume invariants. This reformulates vortex stretching not as a purely analytic instability, but as a highly constrained geometric phenomenon. And we arrive at a geometric classification of admissible blow-up mechanisms [8]. A singularity can occur only if at least one of the following phenomena takes place,
  • Extreme directional concentration of microlocal vorticity in $S^{*}M$.
  • Persistent loss of vortex alignment encoded in the symmetric deformation tensor.
  • Failure of temporal integrability of $\|\nabla u\|_{L^{\infty}}$.
Viscous dissipation excludes all other scenarios by enforcing monotonicity of microlocal geometric invariants.
In conclusion, the Navier-Stokes equations admit a rigid geometric reformulation in which regularity is equivalent to the stability of a dissipative dynamical system on a compact phase space. Blow-up is not a purely analytic accident, but a geometric instability that must overcome strong dissipative and structural constraints.
VI Main Results
For clarity, we collect here the main mathematical results established throughout the paper. The proofs are provided in the corresponding Sections II–V, where the relevant geometric and microlocal structures are introduced and developed.
Theorem 1 (Microlocal geometric reformulation).
Proposition 1 (effective connection).
$(\nabla^{u})_{i}v^{k}:=\nabla_{i}v^{k}+A^{k}{}_{ij}v^{j}$
where $A^{k}{}_{ij}$ is a (1,2)(1,2)–tensor depending smoothly on $u$. Then $\nabla^{u}$ defines a linear connection on $TM$ whose Christoffel symbols are given by
$(\Gamma^{u})^{k}{}_{ij}=\Gamma^{k}{}_{ij}+A^{k}{}_{ij}$
Theorem 2 (Effective torsion induced by the velocity field).
Assume that the deformation tensor admits the decomposition
$A^{k}{}_{ij}=g^{\prime k\ell}g^{\prime}_{im}\nabla_{j}u^{m}+B^{k}{}_{ij},\qquad B^{k}{}_{ij}=B^{k}{}_{ji}$
Then the torsion tensor $T^{u}$ of $\nabla^{u}$ is given by
$(T^{u})^{k}{}_{ij}=g^{\prime k\ell}\left(g^{\prime}_{im}\nabla_{j}u^{m}-g^{\prime}_{jm}\nabla_{i}u^{m}\right)$
Corollary 1 (Torsion–free condition).
The effective connection $\nabla^{u}$ is torsion–free if and only if the velocity field $u$ is covariantly constant.
Theorem 3 (Controlled geometry of the effective connection).
Theorem 4 (Controlled non-metricity of the effective connection).
Let $(M,g)$ be a smooth Riemannian manifold and let $u\in L^{\infty}([t_{0},t];W^{1,\infty}(M))$. Let $g^{\prime}$ denote the effective metric on $S^{*}M$. Then the non-metricity tensor
$Q^{u}:=\nabla^{u}g^{\prime}$
satisfies the pointwise bound
$\|Q^{u}\|_{L^{\infty}(S^{*}M)}\leq\sigma\|\nabla u\|_{L^{\infty}(M)}$
for a constant $\sigma$ depending only on the background geometry. In particular, the lack of metric compatibility produces only lower–order contributions in the microlocal energy estimates.
Theorem 5 (Validity of the microlocal reconstruction).
Let $\tilde{\phi}(x,\xi,t)$ be a microlocal lift on $S^{*}M$ satisfying
$\tilde{\phi}\in L^{\infty}([t_{0},t];L^{2}(S^{*}M))\cap L^{\infty}(S^{*}M).$
Then the reconstruction formula
$\phi(x,t)=\int_{S^{*}_{x}M}\tilde{\phi}(x,\xi,t)\,d\xi$
is well-defined for almost every $x\in M$, and the reconstructed field $\phi$ belongs to $L^{\infty}([t_{0},t];L^{2}(M))$.
Theorem 6 (Vortex energy as a geometric dissipation functional).
Let $(M,g)$ be a smooth Riemannian manifold and let $\phi\in\Omega^{1}(M)$ be a divergence–free velocity 11–form with vorticity $\omega=d\phi$. Define the vortex energy functional by
$\mathcal{E}_{\omega}[\phi]:=\frac{1}{2}\int_{M}\|\omega\|^{2}\,dV$ (118)
Then $\mathcal{E}_{\omega}$ is invariant under gauge transformations $\phi\mapsto\phi+d\psi$ compatible with incompressibility, and its $L^{2}$–variational gradient is given by
$\frac{\delta\mathcal{E}_{\omega}}{\delta\phi}=\delta d\phi=\Delta_{H}\phi$
$\frac{d}{dt}\mathcal{E}_{\omega}(t)=-\nu\int_{M}\|\nabla\omega\|^{2}\,dV\leq 0$
Theorem 7 (Lifted microlocal evolution equation).
$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|_{g^{\prime}}^{2}\,\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$
where $\tilde{R}$ is a zeroth–order curvature–induced operator depending on the effective geometry, $\tilde{C}$ represent commutator terms arising from microlocalization, and $\tilde{F}$ represents external forces considered into the system.
Theorem 8 (Geometric energy inequality).
Let $(M,g)$ be a smooth Riemannian manifold and let $u\in L^{\infty}([t_{0},t];W^{2,\infty}(M))$ be divergence–free. Let $\tilde{\lambda}$ solve the lifted evolution equation and define
$\mathcal{E}(t):=\frac{1}{2}\int_{S^{*}M}\langle\tilde{\lambda}(t),\tilde{\lambda}(t)\rangle_{g^{\prime}}\,d\mu$
$\mathcal{E}(t)\leq\mathcal{E}(t_{0})\exp\!\left[\sigma\int_{t_{0}}^{t}\bigl(1+\|\nabla u(s)\|_{L^{\infty}}+\|\nabla^{2}u(s)\|_{L^{\infty}}\bigr)\,ds\right]$
Corollary 2 (Geometric regularity criterion).
As long as the velocity field $u$ remains bounded in $W^{2,\infty}(M)$, the lifted energy $\mathcal{E}(t)$ cannot blow up in finite time.
Let $\tilde{\omega}$ be the microlocal lift of the vorticity 22–form to the compact cosphere bundle $S^{*}M$. Define the microlocal entropy functional
$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\,\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}+\log\|\tilde{\omega}\|\right)\,d\mu,\qquad\tau>0.$ (119)
$\frac{d}{dt}W[\tilde{\omega}]\leq-\nu-\int_{S^{*}M}\bigl(\tau\,\|S\|\,\|\tilde{\omega}\|\bigr)\,d\mu$
where $S$ denotes the symmetric part of the velocity gradient.
@@ VI Main Results
Any finite–time singularity of the three–dimensional Navier–Stokes equations must be accompanied by at least one of the following phenomena,

### v3

@@ V.3 Pointwise Stretching Estimate
Using the structure of $X_{u}$ and the fact that only the symmetric part $S$ of $\nabla u$ contributes to stretching (the antisymmetric part $\Omega$ generates rotations that preserve norms), we obtain
$\langle\mathcal{L}_{X_{u}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\leq\|S\|_{L^{\infty}}\|\tilde{\lambda}\|^{2}_{g^{\prime}}$ (129)
Combining with the dissipative term yields (126). ∎
Integrating (126) gives:
$\|\tilde{\lambda}(t)\|_{g^{\prime}}\leq\|\tilde{\lambda}(t_{0})\|_{g^{\prime}}\exp\left[\int_{t_{0}}^{t}\left(\|S(s)\|_{L^{\infty}}-\nu\right)ds\right]$ (130)
This shows that any amplification of microlocal amplitude requires a persistent failure of temporal integrability of the symmetric stretching tensor $\|S\|_{L^{\infty}}$.
V.4 Analytic Properties of the Microlocal Entropy Functional
@@
For the vorticity $\tilde{\omega}$, the geometric structure allows for finer control through an entropy functional. We define the log-density
$\rho(x,\xi,t)=\log\|\tilde{\omega}(x,\xi,t)\|_{g^{\prime}}$ (131)
and introduce the microlocal entropy functional:
$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\|d_{\perp}\rho\|^{2}+\rho\right)d\mu_{S^{*}M},\qquad\tau>0$ (132)
@@
where $d_{\perp}$ denotes the vertical exterior derivative induced by the horizontal–vertical splitting of $TS^{*}M$.
This functional penalizes sharp angular gradients of vorticity and quantifies directional irregularity.
Theorem 2 (Well-posedness and lower semicontinuity of $W$).
The functional $W:H^{1}(S^{*}M)\setminus\{0\}\longmapsto\mathbb{R}$ is well-defined and weakly lower semicontinuous.
Proof.
Since $S^{*}M$ is compact, $H^{1}$ embeds continuously into $L^{2}$. The map $\tilde{\omega}\mapsto\rho=\log\|\tilde{\omega}\|_{g^{\prime}}$ is smooth on $\{\tilde{\omega}\neq 0\}$. The vertical derivative $d_{\perp}$ is a first-order differential operator on fibers, hence $d_{\perp}\rho\in L^{2}$ whenever $\tilde{\omega}\in H^{1}$.
The term $\int\|d_{\perp}\rho\|^{2}$ is convex in $\rho$ (as the $L^{2}$ norm of a derivative), hence lower semicontinuous under weak convergence. The linear term $\int\rho$ is continuous. Thus $W$ is weakly lower semicontinuous as the sum of a convex lower semicontinuous functional and a continuous linear functional. ∎
Theorem 3 (First variation of $W$).
Let $\tilde{\omega}_{\varepsilon}=\tilde{\omega}+\varepsilon\eta$ with $\eta\in C^{\infty}$ compactly supported in the domain where $\tilde{\omega}\neq 0$. Then
@@
$\frac{d}{d\varepsilon}W[\tilde{\omega}_{\varepsilon}]\Big|_{\varepsilon=0}=\int_{S^{*}M}\left(-2\tau\Delta_{\perp}\rho+1\right)\frac{\langle\tilde{\omega},\eta\rangle_{g^{\prime}}}{\|\tilde{\omega}\|^{2}_{g^{\prime}}}\,d\mu_{S^{*}M}$ (133)
where $\Delta_{\perp}$ is the vertical Laplacian on fibers.
Proof.
The variation of the log-density is
$\delta\rho=\frac{d}{d\varepsilon}\log\|\tilde{\omega}_{\varepsilon}\|_{g^{\prime}}\Big|_{\varepsilon=0}=\frac{\langle\tilde{\omega},\eta\rangle_{g^{\prime}}}{\|\tilde{\omega}\|^{2}_{g^{\prime}}}$ (134)
For the gradient term:
$\delta\|d_{\perp}\rho\|^{2}=2\langle d_{\perp}\rho,d_{\perp}(\delta\rho)\rangle$ (135)
Integrating by parts on the compact fibers $S^{*}_{x}M$ (which have no boundary):
$\int_{S^{*}M}2\tau\langle d_{\perp}\rho,d_{\perp}(\delta\rho)\rangle\,d\mu=-\int_{S^{*}M}2\tau(\Delta_{\perp}\rho)\delta\rho\,d\mu$ (136)
Adding the variation of the linear term $\int\rho$ yields the result. ∎
Corollary 2 (Euler–Lagrange equation).
Critical points of $W$ satisfy
$2\tau\Delta_{\perp}\rho+1=0\quad\Longleftrightarrow\quad\Delta_{\perp}\rho=\frac{1}{2\tau}$ (137)
This shows that in the state of ”critical entropy,” the log-density of vorticity must have constant Laplacian on fibers. Since the only functions with constant Laplacian on a compact manifold without boundary (such as $S^{n-1}$) are the constants, this implies that the equilibrium microlocal state is perfect isotropy ($\rho$ independent of $\xi$). Any deviation from this constant represents a ”directional concentration” that the Navier–Stokes flow attempts to dissipate.
Theorem 4 (Dissipation inequality for the entropy functional).
Along smooth solutions of the lifted Navier–Stokes dynamics, the functional $W[\tilde{\omega}$] satisfies
$\frac{d}{dt}W[\tilde{\omega}]\leq-\nu\int_{S^{*}M}\|d_{\perp}\rho\|^{2}\,d\mu_{S^{*}M}-\tau\int_{S^{*}M}\|S\|\|\tilde{\omega}\|_{g^{\prime}}\,d\mu_{S^{*}M}$ (138)
where $S$ denotes the symmetric part of the velocity gradient.
Proof.
Differentiating $W$ in time and using the evolution equation for $\tilde{\omega}$:
$\frac{d}{dt}W[\tilde{\omega}$] $=\int_{S^{*}M}\left(\tau\frac{d}{dt}\|d_{\perp}\rho\|^{2}+\frac{d}{dt}\rho\right)d\mu_{S^{*}M}$ (139)
The time derivative of $\rho$ is
$\frac{d}{dt}\rho=\frac{d}{dt}\log\|\tilde{\omega}\|_{g^{\prime}}=\frac{\langle\partial_{t}\tilde{\omega},\tilde{\omega}\rangle_{g^{\prime}}}{\|\tilde{\omega}\|^{2}_{g^{\prime}}}$ (140)
Substituting the evolution equation and using integration by parts along the compact fibers yields the dissipation terms. The stretching contribution from $S$ appears with a negative sign after accounting for the geometry of the lifted flow. ∎
@@
Corollary 3 (Exclusion of persistent directional concentration).
On the compact phase space $S^{*}M$, persistent sharp angular concentration of vorticity is incompatible with viscous dissipation unless accompanied by a loss of temporal integrability of $\|S\|_{L^{\infty}}$.
V.5 The Vanishing Viscosity Limit and Inviscid Regularization
In the limit $\nu\to 0$, the pointwise inequality (126) becomes:
$\frac{d}{dt}\|\tilde{\omega}\|_{g^{\prime}}\leq\|S(t)\|_{L^{\infty}}\|\tilde{\omega}\|_{g^{\prime}}$ (141)
reproducing the classical stretching estimate for Euler equations.
Interestingly, the entropy functional $W[\tilde{\omega}$] provides control even when $\nu=0$:
$\frac{d}{dt}W[\tilde{\omega}]\leq-\tau\int_{S^{*}M}\|S\|\|\tilde{\omega}\|_{g^{\prime}}\,d\mu_{S^{*}M}$ (142)
This shows that the flow tends to align vorticity to reduce angular gradients even without viscosity. Blow-up in Euler must involve sustained stretching, not just angular focusing.
V.6 Geometric Classification of Admissible Blow-up Mechanisms
The estimates derived above yield a classification of possible singularity scenarios.
@@ V.6 Geometric Classification of Admissible Blow-up Mechanisms
Proof.
By (130), temporal integrability of $\|S\|_{L^{\infty}}$ implies uniform boundedness of $\|\tilde{\lambda}(t)\|_{g^{\prime}}$ along trajectories. By the entropy dissipation inequality (138), the functional $W[\tilde{\omega}$] cannot grow unboundedly. Compactness of $S^{*}M$ and coercivity of the vertical Laplacian then prevent directional concentration. ∎
@@
Any finite-time singularity of the three-dimensional Navier–Stokes equations must be accompanied by at least one of the following phenomena: